Circular No. 7949

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 O7 (LINEAR) An apparently asteroidal object reported by LINEAR on July 29.2

UT (discovery observation given below) has been found to be cometary by A. Boattini on CCD images taken with the 0.60-m Schmidt at Campo Imperatore on Aug. 2.9 and by J. Tichá and M. Tichý with the 0.57-m telescope at Kleť on Aug. 4.9 (comet slightly diffuse with a coma of diameter 8′′).

2002 UT  α_{2000}  δ_{2000}  m_{2}
July 29.17312  14^{h}25^{m}04.^{s}75  +46^{o}29′25′′.2  20.0

The available astrometry and the following quite uncertain parabolic orbital elements appear on MPEC 2002-P13:

T = 2003 Sept. 9.767 TT  ω = 254.^{o}585
Ω = 12.413 }2000.0
q = 0.80481 AU  i = 99.088

COMET P/2002 O8 (NEAT) K. Lawrence, Jet Propulsion Laboratory, reports the discovery of a

comet (with nuclear condensation of diameter ∼ 6′′ and a tail extending ∼ 10′′ to the west) on NEAT images obtained on July 29.3 UT (discovery observation appears below). Confirmation of cometary activity was reported by A. C. Gilmore and P. M. Kilmartin on CCD images taken with the 0.6-m reflector at Mt John University Observatory on Aug. 4.4 (faint tail 40′′ in p.a. 255^{o}). Prediscovery observations have been identified in NEAT data from July 16.5, LINEAR data from July 5.4 and Spacewatch data from 2001 May 22.3 and June 11.2.

2002 UT  α_{2000}  δ_{2000}  m_{1}
July 29.32442  21^{h}31^{m}32.^{s}67  −31^{o}01′25′′.2  17.5

The available astrometry and the following orbital elements appear on MPEC 2002-P14:

Epoch = 2002 May 6.0 TT

T = 2002 May 8.5788 TT  ω = 222.^{o}2742
e = 0.199307  Ω = 75.5018 }2000.0
q = 3.229439 AU  i = 12.7858
a = 4.033306 AU n^{o} = 0.1216781 P = 8.100 years

2002 August 5 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother